Assume that the dd & ss function of good Y is given as follows respectively; Qdy =30 - 2Py and Qsy = 16Py - 60. Answer the following questions on the basis of the given information. A) Find the equilibrium price and equilibrium quantity B) What price would bring for the supply of 60 units of good Y? State the condition of the market for the good at the price. C) At what price would no amount of the good be supplied? D) What would be the amount demanded if good Y were a free good?


1
Expert's answer
2022-01-12T08:34:19-0500

A) The equilibrium price and equilibrium quantity are:

Qdy = Qsy,

30 - 2Py = 16Py - 60,

18Py = 90,

Py = 5,

Qy = 30 - 2×5 = 20 units.

B) If Qs = 60 units of good Y, then:

16Py - 60 = 60,

Py = 7.5, so there will be a surplus in this market.

C) No amount of the good be supplied at price of:

16Py - 60 = 0,

Py = 3.75.

D) If good Y were a free good, then the amount demanded is:

Qdy = 30 - 2×0 = 30 units.